# VLBI astrometry of radio stars to link radio and optical celestial reference frames III: 11 radio stars

### v1

@@ Abstract
The alignment between the radio-based International Celestial Reference Frame (ICRF) and the optical Gaia Celestial Reference Frame (Gaia-CRF) is critical for multi-waveband astronomy, yet systematic offsets at the optical bright end () limit their consistency. While radio stars offer a potential link between these frames, their utility has been restricted by the scarcity of precise Very Long Baseline Interferometry (VLBI) astrometry. In this study, we present new VLBI astrometry of 11 radio stars using the Very Long Baseline Array (VLBA), expanding the existing sample with positions, parallaxes, and proper motions measured. All 11 radio stars were detected, for 10 of which parallaxes and proper motions can be estimated, reaching a precision level of in the best cases. These new samples greatly contribute to the link between ICRF and Gaia-CRF at the optical bright end.
@@
The establishment of a unified celestial reference frame (CRF) covering all bands is critical for advancing multi-waveband astronomy, geodesy, and deep-space navigation. At radio wavelengths, the International Celestial Reference Frame (ICRF, 1998AJ....116..516M) is defined through Very Long Baseline Interferometry (VLBI) observations of extragalactic sources (quasars), with its latest iteration, ICRF3, achieving a coordinate noise floor of 30 as (2020A&A...644A.159C). In the optical band, the third data release of the Gaia mission (Gaia DR3, 2023A&A...674A...1G) provides astrometry for over 1.6 million quasars, defining the Gaia Celestial Reference Frame 3 (Gaia-CRF3) with a comparable precision (2022A&A...667A.148G). Gaia-CRF3 is also taken as the optical realization of ICRF3, but in this paper, ICRF3 refers only to its radio realization for simplicity.
@@
Gaia-CRF3 is aligned to ICRF3 at a 7 as level (formal uncertainty) using common quasars (2022A&A...667A.148G). However, systematic discrepancies persist for optical bright sources () in Gaia DR3, attributed to magnitude-, color-, and galactic-latitude-dependent shifts in centroid determinations and instrument calibration limitations, such as the Window Class (WC) effects (2020A&A...633A...1L). These systematics manifest as orientation offsets and residual spins between the bright Gaia-CRF3 and the quasar-based ICRF3 (2018A&A...616A...2L; 2018ApJS..239...31B), necessitating independent validation to achieve an accurate and stable alignment between the CRFs.
@@
Given the optical faintness of most quasars, efforts to link the frames rely on radio stars, which are observable in both radio and optical (bright-end) bands (2016MNRAS.461.1937M). Several studies have attempted to align Gaia-CRF to ICRF with radio stars, e.g., 2020A&A...633A...1L; 2022AstL...48..790B; 2023A&A...676A..11L; 2024A&A...689A.134L; 2024MNRAS.529.2062Z. However, the small number of available radio stars restricts the robustness of the alignment: positions and proper motions of only several tens of radio stars have been measured with VLBI.
@@
It is also possible to assess the bright Gaia-CRF internally within Gaia, which is not restricted by sample numbers. For example, 2021A&A...649A.124C estimated spin parameters with samples collected from known binaries and open clusters. However, orientation parameters cannot be estimated through this method.
@@
It is therefore crucial to increase the number of radio stars with VLBI measurements; to this end, we have undertaken observations within the framework of a long-term research initiative. Observing strategies for radio star astrometry with VLBI are discussed in 2024MNRAS.529.2062Z. Our pilot Very Long Baseline Array (VLBA) program (obs. ID: BZ077) successfully measured parallaxes and proper motion of two radio stars, HD 199178 and AR Lac (2023MNRAS.524.5357C). In another program (VLBA obs. ID: BZ080), we observed three radio stars: HD 179094, RZ Cas, and SZ Psc (Jiang et al. in prep.). This study presents new VLBI astrometry for 11 radio stars, focusing on the data reduction and astrometric results, while the frame link results with all the new data are presented in 2025A&A...699A.345Z.
@@ 2 Observations
Our observations were conducted under VLBA programs BZ087, BZ103, and BZ107. The target selection is introduced in Sect. 2.1, while the observation setups and scheduling are introduced in Sect. 2.2 and 2.3 respectively.
@@ 2.1 Target selection
We selected target radio stars from the Very Large Array (VLA) historical observations. For example, 46 radio stars were observed with the Very Large Array (VLA) Plus Pie Town (2007AJ....133..906B). These stars are detectable on baselines of tens of kilometers; however, it is in doubt whether they are still bright enough to be detected on baselines of thousands of kilometers. Thus, we confirmed through snapshot observation whether they can be detected with VLBI. The snapshot was conducted under the European VLBI Network (EVN) program EZ029 at X band, with 11 antennas (Wb, Ef, Mc, O6, T6, Ur, Ys, Ir, Sv, Zc, Bd), recording rate of 1 Gbps, and on-source time of about five minutes for each star. Six stars (HD 199178, IM Peg, SZ Psc, RZ Cas, RS CVn, and V1762 Cyg) were detected with an SNR 5, in which three (HD 199178, SZ Psc, and RZ Cas) were observed in previous programs (obs. ID: BZ077 and BZ080), and one was selected to be observed in this program (RS CVn).
The Very Large Array Sky Survey (VLASS, 2021ApJS..255...30G) provides a great number of candidates for VLBI observation. Through crossmatching the VLASS catalog with the SIMBAD (simbad) database and Gaia DR2 (2018A&A...616A...1G), we selected 73 radio star candidates that are point-like and have a brightness of 1 mJy/beam. Then we confirmed their detectability through another snapshot observation (EVN obs. ID: EX009) at C band. A recording rate of 2 Gbps, and an on-source time of about ten minutes for each star were adopted, providing twice the sensitivity as EZ029. 20 candidates were detected, and we selected ten to be observed in this program. Information of all 11 selected stars is listed in Table 1, which all have no parallax and proper motion measured with VLBI before. The stars are divided into five groups according to right ascension for easier scheduling.
@@
In the column “Type”, “EB” and “XB” denote Eclipsing Binary and X-ray Binary respectively. Astrometric parameters are from Gaia DR3. and denote position errors in RA(DEC) and DEC directions, respectively. Column “” gives Gaia -band magnitude. RUWE denotes Gaia Renormalized Unit Weight Error.
10 of the 11 selected stars are binaries except for V1859 Ori, with orbital periods ranging from 0.5 to 21 days (2008MNRAS.389.1722E). A reliable estimation of the binary orbits requires an unaffordable amount of observing time for the ten binaries. Fortunately, the angular scales of the binary orbits are one to two orders of magnitude smaller than the value of their parallaxes and proper motions.
@@ 2.2 Observation setups
Most radio stars are weak at radio frequencies (about millijansky level), so we requested the maximum capability of VLBA to achieve a high sensitivity: all 10 antennas, a recording rate of 4 Gbps, dual polarization, and 2-bit quantization. The radiation of radio stars that we can detect mainly has a non-thermal nature, and their fluxes decrease as the frequency increases at centimeter wavelengths (2013tra..book.....W), so C band ( 5 GHz) is a good balance between flux and angular resolution. Ionospheric propagation delay at C band is also smaller than at lower frequencies. On the other hand, the C-band receiver provides the lowest System Equivalent Flux Density (SEFD), i.e., the highest sensitivity, among all VLBA receivers111https://science.nrao.edu/facilities/vlba/docs/manuals/oss/bands-perf.
For weak sources, including radio stars, phase referencing (PR, 1990AJ.....99.1663L; 1995ASPC...82..327B) using a nearby known calibrator enables precise relative astrometry. However, atmospheric spatial structures can introduce additional propagation delays due to the angular separation between the calibrator and the target. To achieve better atmospheric spatial structure correction compared to single-calibrator PR (hereafter referred to simply as PR) and thereby improve astrometric precision, the MultiView technique (2017AJ....153..105R) is used: multiple calibrators that enclose the target are cyclically observed, allowing for a linear phase gradient estimation on the sky plane and interpolation at the target position. The C band is suitable for the application of MultiView, e.g., 2023ApJ...953...21H.
Each target is surrounded by four calibrators, selected from the ICRF3 catalog (2020A&A...644A.159C) and the Radio Fundamental Catalog (RFC, 2025ApJS..276...38P, rfc_2024d used in this work)222https://doi.org/10.25966/dhrk-zh08, which are listed in Table 2. The calibrators are categorized into two types: primary calibrators and secondary calibrators. The primary calibrators serve as absolute position references, meaning the phases of secondary calibrators are residual phases relative to them during the MultiView procedure, which is further introduced in Sect. 3.1. As the primary calibrators are all collected from the ICRF3 catalogue, the absolute positions of the targets we measure are in the ICRF3 system. See Appendix B of 2025AJ....170....4Z for more details about the absolute positioning.
@@ 2.3 Scheduling
Due to the typically low and variable flux densities of radio stars, seven epochs were scheduled for each group of targets to improve the likelihood of successfully measuring their parallaxes and proper motions. The epochs span a period of three years, from 2021 to 2025, instead of the usual one year: BZ087 in semester 2021B, BZ103 in 2024A, and BZ107 in 2024B. This improves the accuracy of proper motions, the most important parameter for CRF spin estimation. The epochs are optimized to sample near the extremes of parallax sinusoids to enable more accurate parallax measurements (2022PASP..134l3001R). The actual epochs for each session are listed in Table 3.
@@ 3.1 Calibration
The raw data were correlated by the VLBA DiFX software correlator (2007PASP..119..318D) in Socorro. A preliminary calibration with standard PR workflow using the primary calibrator was applied to the visibility data:
@@
Manual phase calibration with the AIPS task FRING, i.e., fringe fitting with a fringe finder to remove electronic phase differences among different IF bands.
@@
Most of the non-spatial systematic errors were removed through the PR procedure above. Then serial MultiView (sMV, 2025AJ....170....4Z) was applied to remove the spatial-structure errors. The sMV is an implementation of the MultiView techniques that enables automatic phase ambiguity correction. In order to include the time-domain information and detect “phase wrap” robustly, the sMV operates on the time series of calibrator scans, rather than fitting independent linear phase gradients to each scan group as done in the conventional MultiView implementation. Fringe fitting with the secondary calibrators separately yields their residual phases relative to the primary calibrator, since the fringe fitting result with the primary calibrator has already been applied to them. Under the assumption that the residual phases change continuously, the phase plane should be highly predictable over a short time range, enabling ambiguity detection. The phase plane is iteratively rotated to follow the changes in the residual phase time series, and then the phase correction to be applied to the target can be interpolated in the space and time domain.
We developed a pipeline based on the Astronomical Image Processing System (AIPS, 2003ASSL..285..109G) and ParselTongue (2006ASPC..351..497K), and it is used for all data in this study. For the data used in this work (with the observing sequence “C1-C2-T-C3-C4”), the calibration performance of the sMV pipeline shows virtually no difference from the conventional MultiView implementation, which has been verified in 2025AJ....170....4Z. Details of the whole procedure can also be found there.
After calibration, the targets were imaged and cleaned with the AIPS task IMAGR. An image example is shown in Appendix A. All targets in this study are point-like sources, so their positions were fitted with the AIPS task JMFIT, assuming a single 2-D Gaussian component for each target. All measured positions are available in Appendix B.
@@ 3.2 Astrometric parameter estimation
For stars with 3 detections, position at reference epoch, parallax, and proper motion can be estimated. Here we adopt an astrometric model of five parameters: . Binary orbit, Galactocentric acceleration, and other possible small terms are not taken into consideration. For the parallactic motion, the DE440 (2021AJ....161..105P) ephemeris provided by JPL was used to calculate precise coordinates of the Earth in the Barycentric System. The emcee Markov Chain Monte Carlo sampler (2013PASP..125..306F) was used for parameter estimation.
Usually, errors in VLBI astrometry consist of both thermal error and systematic error (2022PASP..134l3001R). The AIPS task JMFIT calculates a theoretical uncertainty based on the image root mean square (RMS), leading to a significant underestimation of systematic errors. Therefore, systematic error floors in the RA and DEC directions are iteratively added during astrometric parameter estimation to achieve a reduced chi-square . Note this can only be done for stars with 4 detections because the iteration procedure requires an extra degree of freedom.
As discussed in Sect. 2.1, the angular scales of the orbits of the binaries are small. Therefore, a single-star model is used for fitting. And on the other hand, binaries with such short periods and small orbital scales are also fitted with single-star models in the current Gaia astrometric solution; as discussed in 2023A&A...674A...9H, the shortest astrometric binary period ( 30 days) successfully solved in Gaia DR3 is larger than the longest one (AR Mon, 21 days) in our targets.
@@ 3.3 Details of each star
FF Aqr has a slightly variable flux (fluxes mentioned here denote peak flux densities, same below). There is one epoch (BZ103A2) where sMV gives a lower SNR than PR. One of the calibrators, J2156-0333, shows a slight jet-like structure, but it does not introduce significant spurious phases into the residual phase time series. Here, “spurious phase” refers to an evident phase inconsistency between this calibrator and the others on some or all baselines (same definition used below). Self-calibrated images of J2156-0333 and other calibrators with notable structures are shown in Appendix E.

### v2

# myredRGB255,66,56 VLBI astrometry of radio stars to link radio and optical celestial reference frames. II. 11 radio stars

@@ Abstract
The alignment between the radio-based International Celestial Reference Frame (ICRF) and the optical Gaia Celestial Reference Frame (Gaia-CRF) is critical for multi-waveband astronomy, yet systematic offsets at the optical bright end () limit their consistency. While radio stars offer a potential link between these frames, their utility has been restricted by the scarcity of precise Very Long Baseline Interferometry (VLBI) astrometry. In this study, we present new VLBI astrometry of 11 radio stars using the Very Long Baseline Array (VLBA), expanding the existing sample with positions, parallaxes, and proper motions measured. All 11 radio stars were detected, for 10 of which parallaxes and proper motions can be estimated, achieving median uncertainties better than 0.1 mas and 0.1 mas yr-1, respectively. These new samples greatly contribute to the link between ICRF and Gaia-CRF at the optical bright end.
@@
The establishment of a unified celestial reference frame (CRF) covering all bands is critical for advancing multi-waveband astronomy, geodesy, and deep-space navigation. At radio wavelengths, the International Celestial Reference Frame (ICRF, Ma et al., 1998) is the radio realization of the International Celestial Reference System (ICRS, Arias et al., 1995) and is defined through Very Long Baseline Interferometry (VLBI) observations of extragalactic sources (quasars). Its latest iteration, ICRF3, achieved a coordinate noise floor of 30 as (Charlot et al., 2020). In the optical band, the third data release of the Gaia mission (Gaia DR3, Gaia Collaboration et al., 2023) provides astrometry for over 1.6 million quasars, defining the optical realization of the ICRS, the Gaia Celestial Reference Frame 3 (Gaia-CRF3), which has comparable precision to ICRF3 (Gaia Collaboration et al., 2022).
@@
Gaia-CRF3 is aligned to ICRF3 at a 7 as level (formal uncertainty) using common quasars (Gaia Collaboration et al., 2022). However, systematic discrepancies persist for optical bright sources () in Gaia DR3, attributed to magnitude-, color-, and galactic-latitude-dependent shifts in centroid determinations and instrument calibration limitations, such as the Window Class (WC) effects (Lindegren, 2020). These systematics manifest as orientation offsets and residual spins between the bright Gaia-CRF3 and the quasar-based ICRF3 (Lindegren et al., 2018; Brandt, 2018), necessitating independent validation to achieve an accurate and stable alignment between the CRFs.
@@
Given the optical faintness of most quasars, efforts to link the frames rely on radio stars, which are observable in both radio and optical (bright-end) bands (Malkin, 2016). Several studies have attempted to align Gaia-CRF to ICRF with radio stars, e.g., Lindegren (2020); Bobylev (2022); Lunz et al. (2023, 2024); Zhang et al. (2024). However, the small number of available radio stars restricts the robustness of the alignment: positions and proper motions of only several tens of radio stars have been measured with VLBI.
@@
It is also possible to assess the bright Gaia-CRF internally within Gaia, which is not restricted by sample numbers. For example, Cantat-Gaudin and Brandt (2021) estimated spin parameters with samples collected from known binaries and open clusters. However, orientation parameters cannot be estimated through this method.
@@
It is therefore crucial to increase the number of radio stars with VLBI measurements; to this end, we have undertaken observations within the framework of a long-term research initiative. Observing strategies for radio star astrometry with VLBI are discussed in Zhang et al. (2024). Our pilot Very Long Baseline Array (VLBA) program (obs. ID: BZ077) successfully measured parallaxes and proper motion of two radio stars, HD 199178 and AR Lac (Chen et al., 2023). In another program (VLBA obs. ID: BZ080), we observed three radio stars: HD 179094, RZ Cas, and SZ Psc (Jiang et al. in prep.). This study presents new VLBI astrometry for 11 radio stars, focusing on the data reduction and astrometric results, while the frame link results with all the new data are presented in Zhang et al. (2025a).
@@ 2 Observations
Our observations were conducted under VLBA programs BZ087, BZ103, and BZ107. The target selection is introduced in Sect. 2.1, while the observation setups and scheduling are introduced in Sect. 2.2 and 2.3, respectively.
@@ 2.1 Target selection
We selected target radio stars from the Very Large Array (VLA) historical observations. For example, 46 radio stars were observed with the Very Large Array Plus Pie Town (Boboltz et al., 2007). These stars are detectable on baselines of tens of kilometers; however, it is in doubt whether they are bright enough to be detected on baselines of thousands of kilometers. Therefore, we confirmed through snapshot observation whether they can be detected with VLBI. The snapshot was conducted under the European VLBI Network (EVN) program EZ029 at X band, with 11 antennas (Wb, Ef, Mc, O6, T6, Ur, Ys, Ir, Sv, Zc, Bd), recording rate of 1 Gbps, and on-source time of about five minutes for each star. Six stars (HD 199178, IM Peg, SZ Psc, RZ Cas, RS CVn, and V1762 Cyg) were detected with an SNR 5, in which three (HD 199178, SZ Psc, and RZ Cas) were observed in previous programs (obs. ID: BZ077 and BZ080), and one was selected to be observed in this program (RS CVn).
The Very Large Array Sky Survey (VLASS, Gordon et al., 2021) provides a great number of candidates for VLBI observation. Through crossmatching the VLASS catalog with the SIMBAD (Wenger et al., 2000) database and Gaia DR2 (Gaia Collaboration et al., 2018), we selected 73 radio star candidates that are point-like and have a brightness of 1 mJy/beam. Then we confirmed their detectability through another snapshot observation (EVN obs. ID: EX009) at C band. A recording rate of 2 Gbps, and an on-source time of about ten minutes for each star were adopted, providing twice the sensitivity as EZ029. 20 candidates were detected, and we selected ten to be observed in this program. Information of all 11 selected stars is listed in Table 1, which all have no parallax and proper motion measured with VLBI before. The stars are divided into five groups according to their right ascension for easier scheduling.
@@
In the column “Type”, “EB” and “XB” denote Eclipsing Binary and X-ray Binary, respectively. Astrometric parameters are from Gaia DR3. and denote position errors in the RA () and DEC directions, respectively. Column “” gives Gaia -band magnitude. RUWE denotes Gaia Renormalized Unit Weight Error.
10 of the 11 selected stars are binaries except for V1859 Ori, with orbital periods ranging from 0.5 to 21 days (Eker et al., 2008). A reliable estimation of the binary orbits requires an unaffordable amount of observing time for the ten binaries. Fortunately, the angular scales of the binary orbits are one to two orders of magnitude smaller than the value of their parallaxes and proper motions.
@@ 2.2 Observation setups
Most radio stars are weak at radio frequencies (about millijansky level), so we requested the maximum capability of VLBA to achieve a high sensitivity: all 10 antennas, a recording rate of 4 Gbps, dual polarization, and 2-bit quantization. The radiation of radio stars that we can detect mainly has a non-thermal nature, and their fluxes decrease as the frequency increases at centimeter wavelengths (Wilson et al., 2013), so C band ( 5 GHz) is a good balance between flux and angular resolution. Ionospheric propagation delay at C band is also smaller than at lower frequencies. On the other hand, the C-band receiver provides the lowest System Equivalent Flux Density (SEFD), i.e., the highest sensitivity, among all VLBA receivers111https://science.nrao.edu/facilities/vlba/docs/manuals/oss/bands-perf.
For weak sources, including radio stars, phase referencing (PR, Lestrade et al., 1990; Beasley and Conway, 1995) using a nearby known calibrator enables precise differential astrometry. However, atmospheric spatial structures can introduce additional propagation delays due to the angular separation between the calibrator and the target. To achieve better atmospheric spatial structure correction compared to single-calibrator PR (hereafter referred to simply as PR) and thereby improve astrometric precision, the MultiView technique (Rioja et al., 2017) is used: multiple calibrators that enclose the target are cyclically observed, allowing for a linear phase gradient estimation on the sky plane and interpolation at the target position. The C band is suitable for the application of MultiView, e.g., Hyland et al. (2023).
Each target is surrounded by four calibrators, selected from the ICRF3 catalog (Charlot et al., 2020) and the Radio Fundamental Catalog (RFC, Petrov and Kovalev, 2025, rfc_2024d used in this work)222https://doi.org/10.25966/dhrk-zh08, which are listed in Table 2. The calibrators are categorized into two types: primary calibrators and secondary calibrators. The primary calibrators serve as absolute position references, meaning the phases of secondary calibrators are residual phases relative to them during the MultiView procedure, which is further introduced in Sect. 3.1. As the primary calibrators are all collected from the ICRF3 catalogue, the absolute positions of the targets we measure are in the ICRF3 system. See Appendix B of Zhang et al. (2025b) for more details about the absolute positioning.
@@ 2.3 Scheduling
Due to the typically low and variable fluxes of radio stars, seven epochs were scheduled for each group of targets to improve the likelihood of successfully measuring their parallaxes and proper motions. The epochs span a period of three years, from 2021 to 2025, instead of the usual one year: BZ087 in semester 2021B, BZ103 in 2024A, and BZ107 in 2024B. This improves the accuracy of proper motions, the most important parameter for CRF spin estimation. The epochs are optimized to sample near the extremes of parallax sinusoids to enable more accurate parallax measurements (Reid, 2022). The actual epochs for each session are listed in Table 3.
@@ 3.1 Calibration
The raw data were correlated by the VLBA DiFX software correlator (Deller et al., 2007) in Socorro. A preliminary calibration with standard PR workflow using the primary calibrator was applied to the visibility data:
@@
Manual phase calibration with the AIPS task FRING, i.e., fringe fitting with a fringe finder to remove electronic phase differences among different intermediate frequency (IF) bands.
@@
Most of the non-spatial systematic errors were removed through the PR procedure above. Then serial MultiView (sMV, Zhang et al., 2025b) was applied to remove the residual spatial-structure errors. The sMV is an implementation of the MultiView techniques that enables automatic phase ambiguity correction. In order to include the time-domain information and detect “phase wrap” robustly, the sMV operates on the time series of calibrator scans, rather than fitting independent linear phase gradients to each scan group as done in the conventional MultiView implementation. Fringe fitting with the secondary calibrators separately yields their residual phases relative to the primary calibrator, since the fringe fitting result with the primary calibrator has already been applied to them. Under the assumption that the residual phases change continuously, the phase plane should be highly predictable over a short time range, enabling ambiguity detection. The phase plane is iteratively rotated to follow the changes in the residual phase time series, and then the phase correction to be applied to the target can be interpolated in the space and time domain.
We developed a pipeline based on the Astronomical Image Processing System (AIPS, Greisen, 2003) and ParselTongue (Kettenis et al., 2006), and it is used for all data in this study. For the data used in this work (with the observing sequence “C1-C2-T-C3-C4”), the calibration performance of the sMV pipeline shows virtually no difference from the conventional MultiView implementation, which has been verified in Zhang et al. (2025b). Details of the whole procedure can also be found there.
After calibration, the targets were mapped and cleaned with the AIPS task IMAGR. An image example is shown in Appendix A. All targets in this study are point-like sources, so their positions were fitted with the AIPS task JMFIT, assuming a single 2-D Gaussian component for each target. All measured positions are available in Appendix B.
@@ 3.2 Astrometric parameter estimation
For stars with 3 detections, position at reference epoch, parallax, and proper motion can be estimated. Here we adopt an astrometric model of five parameters: . Binary orbit, Galactocentric acceleration, and other possible small terms are not taken into consideration. For the parallactic motion, the DE440 (Park et al., 2021) ephemeris provided by JPL was used to calculate precise coordinates of the Earth in the Barycentric System. The emcee Markov Chain Monte Carlo sampler (Foreman-Mackey et al., 2013) was used for parameter estimation.
Usually, errors in VLBI astrometry consist of both thermal and systematic errors (Reid, 2022). The AIPS task JMFIT calculates a theoretical uncertainty based on the image root mean square (RMS), leading to a significant underestimation of systematic errors. Therefore, systematic error floors in the RA and DEC directions are iteratively added during astrometric parameter estimation to achieve a reduced chi-square . Note this can only be done for stars with 4 detections because the iteration procedure requires an extra degree of freedom in each direction.
As discussed in Sect. 2.1, the angular scales of the orbits of the binaries are small. Therefore, a single-star model is used for fitting. And on the other hand, binaries with such short periods and small orbital scales are also fitted with single-star models in the current Gaia astrometric solution; as discussed in Halbwachs et al. (2023), the shortest astrometric binary period ( 30 days) successfully solved in Gaia DR3 is larger than the longest one (AR Mon, 21 days) in our targets.
@@ 3.3 Details of each star
FF Aqr has a slightly variable flux (the measured values for the point-like radio stars are the peak flux density , same below). There is one epoch (BZ103A2) where sMV gives a lower SNR than PR. One of the calibrators, J2156-0333, shows a slight jet-like structure, but it does not introduce significant spurious phases into the residual phase time series. Here, “spurious phase” refers to an evident phase inconsistency between this calibrator and the others on some or all baselines (same definition used below). Self-calibrated images of J2156-0333 and other calibrators with notable structures are shown in Appendix E.